# Gravitational Waves on Kerr Black Holes I: Reconstruction of Linearized Metric Perturbations

### v2

@@ 1 Introduction and summary
where $\epsilon_{g}=+1$ if the metric has signature (,+,+,+)(-,+,+,+)( - , + , + , + ), or $\epsilon_{g}=-1$ if its signature is (+,,,)(+,-,-,-)( + , - , - , - ). Throughout this paper, we will keep $\epsilon_{g}$ arbitrary to accommodate either choice of convention.
@@
[…]
@@
In keeping with tradition, we will use $\Psi_{\rm H}$ to denote the solution $\Psi^{(-2)}$ to the master equation (1) with $s=-2$ that provides a Hertz potential for $h_{\mu\nu}^{\rm IRG}$, and will use $\Psi_{\rm H}^{\prime}=\zeta^{4}\Psi^{(+2)}$ to denote the (rescaled) solution $\Psi^{(+2)}$ to the master equation (1) with $s=+2$ that provides a Hertz potential for $h_{\mu\nu}^{\rm ORG}$. The Weyl scalars $\psi_{0}$ and $\psi_{4}$ are related to the IRG Hertz potential $\Psi_{\rm H}$ by
@@ 1.2 Structure of the paper
Lastly, we summarize the theory of the confluent Heun function, and carefully derive the quantization conditions for its eigenvalues, in App. E. We also discuss the orthogonality of the confluent Heun functions and derive an analytic formula for their normalization, following the work of Becker [34] on the full Heun equation. Finally, we apply these results to the angular ODE (1.11) in App. F and to the radial ODE (1.10) in App. G. Beyond this paper, we also provide Mathematica notebooks containing our formulas and their checks on this Github.
@@ 2.2 Angular modes
Thus, we can index the discrete set of regular solutions by $(\omega,\ell,m)$, with $\omega\in\mathbb{C}$, $\ell\geq|s|$, and $-\ell\leq m\leq\ell$ [^13] Since the number of zeros $n$ must be nonnegative, $\ell\geq\max\mathopen{}\mathclose{{}\left(|s|,|m|}\right)$, which implies both $\ell\geq|s|$ and $\ell\geq|m|$. These solutions are the eigenmodes $S_{\omega\ell m}^{(s)}(\theta)$ of the angular ODE (1.11), with associated eigenvalues $\lambda_{\omega\ell m}^{(s)}$. These “spin-weighted spheroidal eigenfunctions” take the form
@@
The derivation of these modes is presented below in App. F, which also describes a method for determining the eigenvalues $\lambda_{\omega\ell m}^{(s)}$ from the confluent Heun function. Unfortunately, the $\lambda_{\omega\ell m}^{(s)}$ do not admit a convenient analytical representation, and must always be solved for numerically. The “Black Hole Perturbation Toolkit” BHPToolKit [31] automatically returns them via the command SpinWeightedSpheroidalEigenvalue$[s,\ell,m,a\omega]$.
@@
where the constants $I_{\omega\ell m}^{(s)}$ can be computed numerically, but also admit a closed-form expression in terms of the confluent Heun function and its derivatives, given in App. F.3 below.
@@ 2.3 Radial modes
For every choice of (continuous) frequency $\omega\in\mathbb{C}$ and integer harmonics $(\ell,m)$ with $\ell\geq|s|$ and $-\ell\leq m\leq\ell$, the second-order radial ODE (1.10) admits a two-dimensional space of solutions. A convenient choice of basis for this space is provided by the “in” and “out” solutions, which correspond to modes that are purely ingoing or purely outgoing at the horizon, respectively: [^17] Other common modes are the “up” and “down” solutions, which are purely ingoing or outgoing at infinity (see, e.g., Fig. 1 in [28]). They correspond to solutions of the confluent Heun equation (E.1) around its irregular singular point $z=\infty$, which we unfortunately do not know how to represent in terms of the function $\operatorname{HeunC}(z)$.
@@
The derivation of these modes is presented below in App. G. Their behavior near the horizon is best described using the tortoise coordinate (2.2), which remains regular there:
@@ 2.5 Radial and angular Teukolsky–Starobinsky identities
These operators are the “mode versions” of $l$, $\frac{\Sigma}{\Delta}n$, $\mathcal{L}_{n}$, and ¯$\overline{\mathcal{L}}_{n}$ of Eqs. (1.7) and (1.24), meaning
@@
¯$\overline{\mathcal{L}}_{n}\mathopen{}\mathclose{{}\left[f(r,\theta)e^{-i\omega t+im\phi}}\right]$ $=\mathscr{L}_{n}^{\dagger}f(r,\theta)e^{-i\omega t+im\phi}.$ (2.33d)
@@
$\Delta^{2}\mathopen{}\mathclose{{}\left(\mathscr{D}_{0}^{\dagger}}\right)^{4}\Delta^{2}R_{\omega\ell m}^{(+2)}$ $\propto\hat{R}_{\omega\ell m}^{(-2)}.$ (2.36b)
These statements follow from the assumption that $R_{\omega\ell m}^{(\pm 2)}$ and $S_{\omega\ell m}^{(\pm 2)}$ obey the radial and angular ODEs (1.10) and (1.11), and hold regardless of the linear combination of modes being considered or their normalization. However, if one considers specific modes with a particular normalization, then the proportionality factors become fixed. For the angular modes $\hat{S}_{\omega\ell m}^{(\pm 2)}$ defined in Eq. (2.6),
@@ 2.7 Metric reconstruction in ingoing radiation gauge
$A^{\rm IRG}=0,\qquad B^{\rm IRG}=\frac{4}{\hat{\mathscr{C}}_{\omega\ell m}},$ (2.45)
@@ 3.1 Review of black hole perturbation theory
  • In his Eq. (2-1), the $1$ on the right-hand side becomes $-\epsilon_{g}$, consistent with our Eq. (1.5). (There is also a small typo, as his second inner product should be between $m$ and $\overline{m}$.) Likewise, his expression for the metric as an outer product of the tetrad vectors in Eq. (2-3) must gain a factor of $-\epsilon_{g}$ on its right-hand side.
  • The Newman–Penrose spin coefficients in his Eqs. (2-6) and (2-7) all gain a factor of $-\epsilon_{g}$, so we have, for example,
    $\rho\equiv-\epsilon_{g}\overline{m}^{\mu}m^{\nu}\nabla_{\mu}l_{\nu},\qquad\text{etc.}$. (3.1)
  • The five Weyl scalars in his Eq. (2-11) gain a factor of $-\epsilon_{g}$, so we have, for example,
    $\Psi_{0}\equiv\epsilon_{g}C_{\mu\nu\rho\sigma}l^{\mu}m^{\nu}l^{\rho}m^{\sigma},\qquad\text{etc.}$. (3.2)
@@
  • Due to the factor of $-\epsilon_{g}$ in our definition (1.4) of the Weyl scalars, Price’s Eqs. (3-27) and (3-28) for $\psi_{0}$ and $\psi_{4}$ in terms of the metric perturbation $h_{\mu\nu}$ must gain a factor of $-\epsilon_{g}$.
  • The normalization of the Hertz potentials is not fixed a priori but is set by their relation to the metric components $h_{\mu\nu}$. In our Eqs. (1.18) and (1.20), we included a factor of $2$ (to match Pound and Wardell’s formulas, as aforementioned) as well as a factor of $\epsilon_{g}$, which we inserted so that the relations (1.22) and (1.23) between our Weyl scalars and our Hertz potentials carry no factors of $\epsilon_{g}$. Price chose a different normalization, so his Eq. (3-35) must have its right-hand side multiplied by $-\frac{1}{2}\epsilon_{g}$ in order to recover ours.
@@
    where we have also multiplied by $-\epsilon_{g}$, as explained above.
@@
þ$\text{\thorn}^{4}\Psi_{2}^{-4/3}\psi_{4}$ =ð$={\eth^{\prime}}^{4}\Psi_{2}^{-4/3}\psi_{0}-3\mathcal{V}\overline{\psi}_{0},$ (3.7a)
þ${\text{\thorn}^{\prime}}^{4}\Psi_{2}^{-4/3}\psi_{0}$ =ð$=\eth^{4}\Psi_{2}^{-4/3}\psi_{4}+3\mathcal{V}\overline{\psi}_{4}.$ (3.7b)
@@ 3.4 Metric reconstruction
Finally, we can put everything together to derive the metric reconstruction formulas presented in Secs. 2.7 and 2.8. The metric perturbations $h_{\mu\nu}$ in IRG and ORG are given by Eqs. (1.18) and (1.20), respectively, with the operators ${S_{0}^{\dagger}}_{\mu\nu}$ and ${S_{4}^{\dagger}}_{\mu\nu}$ taking the forms given in Eqs. (3.10). We apply the operators in this form to the mode expansions (3.11) of the IRG and ORG Hertz potentials. Since each operator has three terms, we see that the metric components $h_{\mu\nu}$ can all be expressed in terms of three functions, which we called $H_{\omega\ell m}^{ab}(r,\theta)$ in Secs. 2.7 and 2.8. The computation of these functions is essentially the main nontrivial step in metric reconstruction.21
As an example, we compute $H_{\omega\ell m}^{nm}$ in the ingoing radiation gauge. This requires computing $h_{nm}^{\rm IRG}$ by first acting with the form (3.10) of ${S_{0}^{\dagger}}_{\mu\nu}$ on the Hertz potential $\Psi_{\rm H}$, then plugging the result into Eq. (1.18) to obtain $h_{\mu\nu}^{\rm IRG}$, before finally projecting as in Eq. (2.5) to find
@@
Each of these terms takes the form of a second-order differential operator applied to one of the modes of the Hertz potential. We can now replace all the operators $l$, $\frac{\Sigma}{\Delta}n$, $\mathcal{L}_{n}$, and ¯$\overline{\mathcal{L}}_{n}$ by their “mode versions” (2.33), which show how they act on modes behaving like $e^{-i\omega t+im\phi}$. If instead, they act on complex-conjugated modes behaving like $e^{i\overline{\omega}t-im\phi}$, then one must replace $(\omega,m)\to(-\overline{\omega},-m)$ in the operators (2.31) and (2.32), which amounts to sending
@@ Appendix B Metric reconstruction in ingoing and outgoing coordinates
In this appendix, we take the reconstructed metrics in ingoing and outgoing radiation gauge (given in Secs. 2.7 and 2.8, respectively) and transform their Boyer-Lindquist components $h_{\mu\nu}^{\rm IRG/ORG}$ to ingoing coordinates (in App. B.1) and to outgoing coordinates (in App. B.2). We expect these forms of the metric components to be particularly useful for analyses of the peturbations at (or near) the horizon, where the Boyer-Lindquist coordinates become singular, while the ingoing and outgoing coordinates remain regular. Another advantage they provide is that many components of $h_{\mu\nu}^{\rm ORG}$ (or $h_{\mu\nu}^{\rm IRG}$) vanish in ingoing (or resp., outgoing) coordinates.
@@ B.1 Metric reconstruction in ingoing coordinates
$h_{r\psi}^{\rm IRG}$ $=\frac{2\sin{\theta}\Sigma}{\Delta}\mathopen{}\mathclose{{}\left[-\mathopen{}\mathclose{{}\left(r^{2}+a^{2}}\right)\mathcal{N}_{-}+a\sin{\theta}h_{nn}}\right],$ (B.4i)
@@ Appendix C Eliminating derivatives of the mode functions
$\hat{D}_{\omega\ell m}S_{\omega\ell m}^{(-2)}$ $=-\mathbb{X}_{\omega\ell m}\mathscr{L}_{2}\hat{S}_{\omega\ell m}^{(+2)}+\mathbb{Y}_{-\omega,\ell,-m}\hat{S}_{\omega\ell m}^{(+2)},$ (C.1d)
where we (yet again) introduced some new coefficients18
@@ E.1 Canonical form of the confluent Heun equation
$q$ is known as the accessory parameter. The full Heun equation has four regular singular points. This confluent version arises when two of them coalesce, leaving two regular singular points at $z=0$ and $z=1$ and an irregular singular point at $z=\infty$. As a second-order ODE, it admits two independent solutions. The lowest powers of $z$ in the series expansions of the two solutions about $z=0$ are 0 and $1-\gamma$. As for the two series expansions about $z=1$, their lowest powers of $1-z$ are 0 and $1-\delta$. We let $\operatorname{HeunC}(q,\alpha,\gamma,\delta,\epsilon;z)$ denote the unique solution whose power series expansion around $z=0$ has the leading term $z^{0}$ with coefficient 1. This function is hard-coded in Mathematica as HeunC$[q,\alpha,\gamma,\delta,\epsilon,z]$ and is a “special function” of mathematical physics.
In terms of this special function, the local Frobenius solutions around $z=0$ are [^24] When $1-\gamma$ is integer, one of the solutions is replaced by a more complicated expression involving a different power series plus a piece proportional to the other solution multiplied by $\ln(z)$. If $01-\gamma>$ 0, then the leading term in the series is $z^{0}$, and the solution multiplying the logarithm has leading term $z^{1-\gamma}$. The opposite holds for $01-\gamma<$ 0. Similar statements can be made for the solutions around $z=1$ when $1-\delta$ is integer.

### v3

@@ 1 Introduction and summary
where $\epsilon_{g}=+1$ if the metric has signature (,+,+,+)(-,+,+,+), or $\epsilon_{g}=-1$ if its signature is (+,,,)(+,-,-,-). Throughout this paper, we will keep $\epsilon_{g}$ arbitrary to accommodate either choice of convention.
@@
The problem of metric reconstruction was first tackled by Chrzanowski [20] and Cohen and Kegeles [21], who found explicit formulas for solutions $h_{\mu\nu}$ of the linearized Einstein equations (1.2) in terms of solutions $\Psi^{(\pm 2)}$ of the Teukolsky master equation (1). Soon after, Wald [22] used a notion of adjointness (a $\dagger$} operation) for tensorial differential operators to give a short and elegant derivation of these results (see App. C of [23] for a concise summary of his approach). Similar equations for constructing spacetime fields associated with spin-$s$ perturbations from $\Psi^{(\pm s)}$ were also derived by Cohen and Kegeles [24] and many other later authors.
@@
In keeping with tradition, we will use $\Psi_{\rm H}$ to denote the solution $\Psi^{(-2)}$ to the master equation (1) with s=2s=-2 that provides a Hertz potential for $h_{\mu\nu}^{\rm IRG}$, and will use $\Psi_{\rm H}^{\prime}=\zeta^{4}\Psi^{(+2)}$ to denote the (rescaled) solution $\Psi^{(+2)}$ to the master equation (1) with $s=+2$ that provides a Hertz potential for $h_{\mu\nu}^{\rm ORG}$. The Weyl scalars $\psi_{0}$ and $\psi_{4}$ are related to the IRG Hertz potential $\Psi_{\rm H}$ by
@@ 1.2 Structure of the paper
Lastly, we summarize the theory of the confluent Heun function, and carefully derive the quantization conditions for its eigenvalues, in App. E. We also discuss the orthogonality of the confluent Heun functions and derive an analytic formula for their normalization, following the work of Becker [34] on the full Heun equation. Finally, we apply these results to the angular ODE (1.11) in App. F and to the radial ODE (1.10) in App. G. Beyond this paper, we also provide Mathematica notebooks containing our formulas and their checks on this Github. Updated link.
@@ 2.2 Angular modes
Thus, we can index the discrete set of regular solutions by $(\omega,\ell,m)$, with $\omega\in\mathbb{C}$, $\ell\geq|s|$, and $\ell\leq m\leq\ell$. [^13] Since the number of zeros $n$ must be nonnegative, $\ell\geq\max\mathopen{}\mathclose{{\left(|s|,|m|}}\right)$, which implies both $\ell\geq|s|$ and $\ell\geq|m|$. These solutions are the eigenmodes $S_{\omega\ell m}^{(s)}(\theta)$ of the angular ODE (1.11), with associated eigenvalues $\lambda_{\omega\ell m}^{(s)}$. These “spin-weighted spheroidal eigenfunctions” take the form
@@
[…]
@@
where the constants $I_{\omega\ell m}^{(s)}$ can be computed numerically [removed comma] but also admit a closed-form expression in terms of the confluent Heun function and its derivatives, given in App. F.3 below.
@@ 2.3 Radial modes
For every choice of (continuous) frequency $\omega\in\mathbb{C}$ and integer harmonics $(\ell,m)$ with $\ell\geq|s|$ and $\ell\leq m\leq\ell$, the second-order radial ODE (1.10) admits a two-dimensional space of solutions. A convenient choice of basis for this space is provided by the “in” and “out” solutions, which correspond to modes that are purely ingoing or purely outgoing at the horizon, respectively: [^17] Other common modes are the “up” and “down” solutions, which are purely ingoing or outgoing at infinity (see, e.g., Fig. 1 in [28]). They correspond to solutions of the confluent Heun equation (E.1) around its irregular singular point $z=\infty$, which we unfortunately do not know how to represent in terms of the function $\operatorname{HeunC}(z)$.
@@
The derivation of these modes is presented below in App. G. Their behavior near the horizon is best described using the tortoise coordinate (2.2):
@@ 2.5 Radial and angular Teukolsky–Starobinsky identities
These operators are the “mode versions” of $l$, $\frac{\Sigma}{\Delta}n$, $\mathcal{L}_{n}$, and $\overline{\mathcal{L}}_{n}$ of Eqs. (1.7) and (1.24), meaning
@@
$\overline{\mathcal{L}}_{n}\mathopen{}\mathclose{{\left[f(r,\theta)e^{-i\omega t+im\phi}}}\right]$ $=\mathscr{L}_{n}^{\dagger}f(r,\theta)e^{-i\omega t+im\phi}.$ (2.33d)
@@
$\Delta^{2}\mathopen{}\mathclose{{\left(\mathscr{D}_{0}^{\dagger}}}\right)^{4}\Delta^{2}R_{\omega\ell m}^{(+2)}$ $\propto R_{\omega\ell m}^{(-2)}.$ (2.36b)
These statements follow from the assumption that $R_{\omega\ell m}^{(\pm 2)}$ and $S_{\omega\ell m}^{(\pm 2)}$ obey the radial and angular ODEs (1.10) and (1.11) and hold regardless of the linear combination of modes being considered or their normalization. However, if one considers specific modes with a particular normalization, then the proportionality factors become fixed. For the angular modes $\hat{S}_{\omega\ell m}^{(\pm 2)}$ defined in Eq. (2.6),
@@ 2.7 Metric reconstruction in ingoing radiation gauge
$A^{\rm IRG}=0,\qquad B^{\rm IRG}=\frac{4}{\overline{\hat{\mathscr{C}}}_{\omega\ell m}},$ (2.45)
@@ 3.1 Review of black hole perturbation theory
  • In his Eq. (2-1), the $1$ on the right-hand side becomes $\epsilon_{g}$, consistent with our Eq. (1.5). (There is also a small typo, as his second inner product should be between $m$ and $\overline{m}$.) Likewise, his expression for the metric as an outer product of the tetrad vectors in Eq. (2-3) must gain a factor of $\epsilon_{g}$ on its right-hand side.
  • The Newman–Penrose spin coefficients in his Eqs. (2-6) and (2-7) all gain a factor of $\epsilon_{g}$, so we have, for example,
    $\rho\equiv-\epsilon_{g}\overline{m}^{\mu}m^{\nu}\nabla_{\mu}l_{\nu},\qquad\text{etc.}$ (3.1)
  • The five Weyl scalars in his Eq. (2-11) gain a factor of $\epsilon_{g}$, so we have, for example,
    $\Psi_{0}\equiv\epsilon_{g}C_{\mu\nu\rho\sigma}l^{\mu}m^{\nu}l^{\rho}m^{\sigma},\qquad\text{etc.}$ (3.2)
@@
  • Due to the factor of $\epsilon_{g}$ in our definition (1.4) of the Weyl scalars, Price’s Eqs. (3-27) and (3-28) for $\psi_{0}$ and $\psi_{4}$ in terms of the metric perturbation $h_{\mu\nu}$ must gain a factor of $\epsilon_{g}$.
  • The normalization of the Hertz potentials is not fixed a priori but is set by their relation to the metric components $h_{\mu\nu}$. In our Eqs. (1.18) and (1.20), we included a factor of $2$ (to match Pound and Wardell’s formulas, as aforementioned) as well as a factor of $\epsilon_{g}$, which we inserted so that the relations (1.22) and (1.23) between our Weyl scalars and our Hertz potentials carry no factors of $\epsilon_{g}$. Price chose a different normalization, so his Eq. (3-35) must have its right-hand side multiplied by $\frac{1}{2}\epsilon_{g}$ in order to recover ours.
@@
    where we have also multiplied by $\epsilon_{g}$, as explained above.
@@
$\text{\thorn}^{4}\Psi_{2}^{-4/3}\psi_{4}$ $={\eth^{\prime}}^{4}\Psi_{2}^{-4/3}\psi_{0}-3\mathcal{V}\overline{\psi}_{0},$ (3.7a)
${\text{\thorn}^{\prime}}^{4}\Psi_{2}^{-4/3}\psi_{0}$ $=\eth^{4}\Psi_{2}^{-4/3}\psi_{4}+3\mathcal{V}\overline{\psi}_{4}.$ (3.7b)
@@ 3.4 Metric reconstruction
Finally, we can put everything together to derive the metric reconstruction formulas presented in Secs. 2.7 and 2.8. The metric perturbations $h_{\mu\nu}$ in IRG and ORG are given by Eqs. (1.18) and (1.20), respectively, with the operators ${\mathcal{S}_{0}^{\dagger}}_{\mu\nu}$ and ${\mathcal{S}_{4}^{\dagger}}_{\mu\nu}$ taking the forms given in Eqs. (3.10). We apply the operators in this form to the mode expansions (3.11) of the IRG and ORG Hertz potentials. Since each operator has three terms, we see that the metric components $h_{\mu\nu}$ can all be expressed in terms of three functions, which we called $H_{\omega\ell m}^{ab}(r,\theta)$ in Secs. 2.7 and 2.8. The computation of these functions is essentially the main nontrivial step in metric reconstruction.21
As an example, we compute $H_{\omega\ell m}^{nm}$ in the ingoing radiation gauge. This requires computing $h_{nm}^{\rm IRG}$ by first acting with the form (3.10) of ${\mathcal{S}_{0}^{\dagger}}_{\mu\nu}$ on the Hertz potential $\Psi_{\rm H}$, then plugging the result into Eq. (1.18) to obtain $h_{\mu\nu}^{\rm IRG}$, before finally projecting as in Eq. (2.5) to find
@@
Each of these terms takes the form of a second-order differential operator applied to one of the modes of the Hertz potential. We can now replace all the operators $l$, $\frac{\Sigma}{\Delta}n$, $\mathcal{L}_{n}$, and $\overline{\mathcal{L}}_{n}$ by their “mode versions” (2.33), which show how they act on modes behaving like $e^{-i\omega t+im\phi}$. If instead, they act on complex-conjugated modes behaving like $e^{i\overline{\omega}t-im\phi}$, then one must replace $(\omega,m)\to(-\overline{\omega},-m)$ in the operators (2.31) and (2.32), which amounts to sending
@@ Appendix B Metric reconstruction in ingoing and outgoing coordinates
In this appendix, we take the reconstructed metrics in ingoing and outgoing radiation gauge (given in Secs. 2.7 and 2.8, respectively) and transform their Boyer-Lindquist components $h_{\mu\nu}^{\rm IRG/ORG}$ to ingoing coordinates (in App. B.1) and to outgoing coordinates (in App. B.2). We expect these forms of the metric components to be particularly useful for analyses of the perturbations at (or near) the horizon, where the Boyer-Lindquist coordinates become singular, while the ingoing and outgoing coordinates remain regular. Another advantage they provide is that many components of $h_{\mu\nu}^{\rm ORG}$ (or $h_{\mu\nu}^{\rm IRG}$) vanish in ingoing (or resp., outgoing) coordinates.
@@ B.1 Metric reconstruction in ingoing coordinates
$h_{r\psi}^{\rm IRG}$ $=\frac{2\Sigma\sin{\theta}}{\Delta}\mathopen{}\mathclose{{\left[-\mathopen{}\mathclose{{\left(r^{2}+a^{2}}}\right)\mathcal{N}_{-}+a\sin{\theta}h_{nn}}}\right],$ (B.4i)
@@ Appendix C Eliminating derivatives of the mode functions
$\hat{D}_{\omega\ell m}\hat{S}_{\omega\ell m}^{(-2)}$ $=-\mathbb{X}_{\omega\ell m}\mathscr{L}_{2}\hat{S}_{\omega\ell m}^{(+2)}+\mathbb{Y}_{-\omega,\ell,-m}\hat{S}_{\omega\ell m}^{(+2)},$ (C.1d)
[Hat missing here.] where we (yet again) introduced some new coefficients18
@@ E.1 Canonical form of the confluent Heun equation
[…]
In terms of this special function, the local Frobenius solutions around $z=0$ are [^24] When $1-\gamma$ is integer, one of the solutions is replaced by a more complicated expression involving a different power series plus a piece proportional to the other solution multiplied by $\ln(z)$. If $1-\gamma>0$, then the leading term in the series is $z^{0}$, and the solution multiplying the logarithm has leading term $z^{1-\gamma}$. The opposite holds for $1-\gamma<0$. Similar statements can be made for the solutions around $z=1$ when $1-\delta$ is integer.